OpenStudy (anonymous):

Choose the correct simplification of x to the 6th power over x to the 3rd power. Please someone help?

OpenStudy (akashdeepdeb):

\[x^6/x^3 = x^3\] Because x^(6-3) = x^3!! Understood? :D

OpenStudy (anonymous):

yes

OpenStudy (akashdeepdeb):

Great! :)

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

can you help me with one more please?

OpenStudy (anonymous):

Choose the correct simplification of (2a2 + 7a − 3) + (4a2 − 3).